OpenStudy (anonymous):

Tracy ran for 5/6 hours and Reina ran for 1/2 hours. How many more hours did Tracy run?

OpenStudy (anonymous):

Get them into a common denominator. It's asking for this: \[\frac 56 - \frac 12 = ?\]

OpenStudy (anonymous):

(5/6)-(1/2) Find the common denominator.

OpenStudy (anonymous):

that would be 6

OpenStudy (anonymous):

If the common denominator is 6, then this is what you do. \[\frac 56 - \frac 12 = \frac56 - (\frac 12 \times \frac33) = \frac 56 - \frac36 = ?\]

OpenStudy (anonymous):

answer is 1/3

OpenStudy (anonymous):

What is 5-3?

OpenStudy (anonymous):

Yes it is :)

OpenStudy (anonymous):

\[\frac 56 - \frac 36 = \frac 26 = \frac 13\]\[\large Good \space Job! \space ;)\]